Two cars start from the same place and travel in opposite directions. One runs at an average rate of 24 kilometers per hour faster than the other. Find their rates if after 6 hours, they are 702 km apart.
Answer by mananth(16946)

let rate of one car be xkph
other car rate will be x+24 kph
..
distance 702 from each other after 6 hours
since they are going away from each other the effective rate will be x+x+24 = 2x+24 kph
..
distance = rate * time
702 = (2x+24) * 6
702/6 = 2x+24
117 =2x+24
117-24 = 2x
93=2x
x= 46.5 kph
other car rate = x+24 = 46.5 +24 = 70.5 kph
...
m.ananth@hotmail.ca